jesternmais:

Robin and Kelly own neighboring cornfields. Robin harvested 4 and 3/10 acres of corn on Monday and Kelly harvested 2 and 1/10 acres. How many more acres did Robin harvest than Kelly?

simsharrison:

by 4 and 3/10 do you mean 4 3/10

simsharrison:

@jhonyy9

TheTharapist:

yea

XTHb:

dont give direct answers

XTHb:

thats against the rules

TheTharapist:

WHAT.

XTHb:

mhm so please refrain from that

jesternmais:

ok

jesternmais:

am a try

simsharrison:

You need to understand how to work with fractions: 4 3/10+2 1/10 = (4+2) ( 3/10+1/10) = 6 4/10 = 6 2/5

simsharrison:

@jhonyy9 ? is that right

jesternmais:

oh I really need to its hard tho I jus stated it again

jesternmais:

but thxx am a take note if it

jhonyy9:

\[4\frac{ 3 }{ 10 } = \frac{ 43 }{ 10 }\] \[2\frac{ 1 }{ 10 } = \frac{ 21 }{ 10 }\] make the difference and will get the right answer

jhonyy9:

ok ?

simsharrison:

@jesternmais do you know how to subtract fractions?

jhonyy9:

\[\frac{ 43 }{ 10 } - \frac{ 21 }{ 10 } = ?\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison You need to understand how to work with fractions: 4 3/10+2 1/10 = (4+2) ( 3/10+1/10) = 6 4/10 = 6 2/5 \(\color{#0cbb34}{\text{End of Quote}}\) yes right just simplified this is why not is the same with my result

simsharrison:

It is, you simplified it better, well done!

jesternmais:

thank you